OpenStudy (anonymous):

solve the equation 2 tan c-3=3 tan C-4 algebraically for all values of c in the interval 0≤C<360

OpenStudy (anonymous):

@celecity plz helppp

OpenStudy (anonymous):

@AccessDenied

OpenStudy (anonymous):

@satellite73

OpenStudy (accessdenied):

Are these the same variable C or is there a difference between lowercase and uppercase c/C's?

OpenStudy (anonymous):

its same?

OpenStudy (accessdenied):

Okay. Have you made any attempts so far? Or not sure how to begin? \( 2 \tan C - 3 = 3 \tan C - 4\) \( 0 \le C < 360\)

OpenStudy (anonymous):

do u add 3 on both sideS?

OpenStudy (accessdenied):

That would be fine. \( 2 \tan C = 3 \tan C - 1\) Then we can move 3tan C to the left side.

OpenStudy (anonymous):

im gona do 2tanc -3=3tanc-4?

OpenStudy (accessdenied):

From 2 tan C = 3 tan C - 1 We can subtract the entire 3 tan C from both sides to group it on the left side. Can you see how that works out? 2 tan C = 3 tan C - 1 2 tan C - 3 tan C = 3 tan C - 1 - 3 tan C (2 - 3) tan C = 0 - 1

OpenStudy (anonymous):

ugh im totally confused on what ur doin can u guide me step by step plz?

OpenStudy (accessdenied):

So, I'll start from the initial equation: \( 2 \tan C - 3 = 3 \tan C - 4 \) You were correct to say to add 3 to both sides. I will colorize the additions made. \( 2 \tan C - 3 \color{green}{ + 3} = 3 \tan C - 4 \color{Green}{+ 3} \) \( 2 \tan C = 3 \tan C - 1 \) This is where we were moments ago. The next step is as if we had this equation with u in place of tan C. \( 2 u = 3 u - 1 \) You would subtract 3u from both sides to solve for u, yes?

OpenStudy (accessdenied):

\( 2u = 3u - 1 \) ==> \( 2u \color{Green}{- 3u} = 3u - 1 \color{green}{- 3u} \) ==> \(-u = -1\) \( 2 \tan C = 3 \tan C - 1 \) ==> \( 2 \tan C \color{green}{- 3 \tan C} = \cancel{3 \tan C} - 1 \color{green}{\cancel{- 3 \tan C}} \) ==> \( (2 - 3) \tan C = -1 \) ==> \( - \tan C = -1 \)

OpenStudy (anonymous):

but y r u choosing u over c?

OpenStudy (accessdenied):

u is replacing tan C, to make a point about how to solve. u = tan C would make this look more like a simple single-variable equation like "2u - 5 = 6"

OpenStudy (accessdenied):

It's not significant to the problem itself, but rather to make it more evident to you that the solution comes from simply trying to solve for the "tan C" term first.

OpenStudy (anonymous):

oh like x? 3x-2=0 lets say

OpenStudy (accessdenied):

Yes. If you wanted to solve for that x, you would take a similar route. It is the same with the entire "tan C" part and u. \( 2x = 3x - 1 \) \( 2\tan C = 3 \tan C - 1 \) \(2x \color{green}{- 3x} = 3x - 1 \color{Green}{- 3x}\) \(2 \tan C \color{green}{- 3 \tan C} = 3 \tan C - 1 \color{green}{- 3 \tan C} \) \((2-3) x = -1 \) \((2 - 3) \tan C = -1 \) \(-1 x = -1\) \(-1 \tan C = -1 \)

OpenStudy (anonymous):

ok gt it now im up to the step wher i got -tanc=-1

OpenStudy (anonymous):

what do i do afta?

OpenStudy (accessdenied):

you'd want "tan C" alone, so the coefficient needs to go to the right side. we'd multiply both sides by -1 (or divide, both are same operation...) \( -x = -1\) \( -\tan C = -1 \) \(-1 * -x = -1 * -1 \) \( -1 * -\tan C = -1 * -1 \)

OpenStudy (accessdenied):

which gets you tan C = 1.

OpenStudy (anonymous):

ok i got tht,.

OpenStudy (anonymous):

den?

OpenStudy (accessdenied):

Last part is finding what values of C will get you a value for tangent of 1. If you know the unit circle, this is one to find on that. tan C = (sin C)/(cos C), so where sin C = cos C. tan C = 1 0 <= C < 360

OpenStudy (accessdenied):

or a calculator may also tell you the value here: tan C = 1 C = tan^(-1) (1) but because it is restricted to -180 to 180, you will miss any values 180< C< 360.

OpenStudy (anonymous):

I got -45?

OpenStudy (accessdenied):

is that tan^(-1) (1) or tan^(-1) (-1) ?

OpenStudy (accessdenied):

by tan^(-1) (1) I mean: \( \tan^{-1} \left(1\right) \)

OpenStudy (anonymous):

i stl gt -45

OpenStudy (accessdenied):

weird. that looks like the value for \( \tan^{-1} \left(-1 \right) \)

OpenStudy (anonymous):

i type \[\tan^(-1) \]

OpenStudy (anonymous):

tan −1 (−1)

OpenStudy (anonymous):

i type 2ndtan(-1)

OpenStudy (accessdenied):

Do you agree we ended up with this equation? tan C = 1 <-- positive one?

OpenStudy (anonymous):

yes i gt tht

OpenStudy (accessdenied):

we take inverse tangent of both sides: \( \tan^{-1} \left( \tan C \right) = \tan^{-1} \left( 1 \right) \) \( C = \tan^{-1} \left(1 \right) \) Apologies if the nonformatted answer was confusing. But you should type 2nd tan(1) for positive 45 degrees.

OpenStudy (anonymous):

ohh now i got 45!

OpenStudy (accessdenied):

yep. :) the only other thing to note is that this was only the solutions in -90 < C < 90, the range of inverse tangent. for 90 < C < 270 , the other solution is 45 + 180 = 225 degrees. this is because tangent has a period of 180 degrees (pi radians).

OpenStudy (accessdenied):

so for C: tan (C + 180) = tan C

OpenStudy (anonymous):

how did u no to do tht tho?

OpenStudy (accessdenied):

we wanted to find all the values of C over this interval: 0 <= C < 360, a given. tangent is a periodic function with period 180. we restrict its domain to [-90, 90] to make the inverse function, a function with a single value. when we took tan ^(-1) (1), we obtained the value for C in the domain of tangent [-90, 90]. but in general, any value: C + 180n will give the same value as C because tangent is periodic. tan (C + 180) = tan C (a) We found C = 45 as one solution. What about 45 + 180 n ? (b) Adding 180 to 45, C = 180+45 = 225 is also in our interval 0 <= C < 360. (c) Adding 180 again to 225, 405 is NOT in our interval of values for C.

OpenStudy (accessdenied):

if you had written C = 45 + 180n from the start (and did this every time), the problem becomes "how many integer values of n will get us a value of C in the interval: 0 <= C < 360 ? " C = 45 + 180(0) = 45 C = 45 + 180(1) = 225

OpenStudy (anonymous):

ohh i get now thanks a lot ur awesome!

OpenStudy (accessdenied):

glad to help! :)

OpenStudy (anonymous):

ru a teacher by any chance?

OpenStudy (accessdenied):

nope, i am a high school senior. :D

OpenStudy (anonymous):

wow nice u had the patience to explain it to me. wt u state do u liv in?

OpenStudy (accessdenied):

Illinois. :) Yes, I have a lot of patience for Math. (I recall spending 6 hours or more helping someone on OS with geometry one night, heh).

OpenStudy (anonymous):

wow thts crzy i have another ques last one for 2night tho lol

OpenStudy (anonymous):

find algebraically, the measure of the obtuse angle to the nearest degree, tht satisfies 5 csc theta=8

OpenStudy (accessdenied):

\( 5 \csc \theta = 8 \) I would recommend converting reciprocal functions like cosecant and secant to sines and cosines. It makes them easier to deal with. \( 5 \dfrac{1}{\sin \theta} = 8 \)

OpenStudy (anonymous):

ok then?

OpenStudy (accessdenied):

Multiply both sides by sin theta. \( 5 = 8 \sin \theta \) and then get rid of coefficient on sin theta by dividing it off both sides: \( \dfrac{5}{8} = \dfrac{\cancel{8}}{\cancel{8}} \sin \theta \) that makes sense so far?

OpenStudy (anonymous):

multiply the 8 by theta? and wht else/?

OpenStudy (accessdenied):

Uhh, we won't need to multiply by anything from here. We have sin theta = a fraction. Now we can take inverse sine of both sides.

OpenStudy (anonymous):

no im stil up to the 1st step

OpenStudy (accessdenied):

Oh, this was a division by theta: \( 5 = 8 \sin \theta \) \( \dfrac{5}{8} = \sin \theta \)

OpenStudy (accessdenied):

ERR, division by eight ***

OpenStudy (anonymous):

how r u gtting 5=8sin theata?

OpenStudy (accessdenied):

oh i'm sorry, i see what you were questioning. \( 5 \csc \theta = 8 \) \( 5 \dfrac{1}{\sin \theta} = 8 \) <-- multiplying both sides by \( \sin \theta \) \( 5 \dfrac{1}{\sin \theta } \times \sin \theta = 8 \times \sin \theta \)

OpenStudy (accessdenied):

the sin theta divided by sin theta cancels on the left, leaving 5. \( 5 = 8 \sin \theta \)

OpenStudy (anonymous):

what happened to the one?

OpenStudy (accessdenied):

\( \dfrac{1}{\sin \theta} \times \sin \theta = \dfrac{1}{\sin \theta} \times \dfrac{\sin \theta}{1} \) dividing by 1 does not change value. You multiply fractions straight across, that sounds familiar?

OpenStudy (anonymous):

right so now im up to 5=8sintheata

OpenStudy (accessdenied):

yup. \( 5 = 8 \sin \theta \) did it make sense when I said solve for \( \sin \theta \) by removing the coefficient? ( like 5 = 8 x )

OpenStudy (anonymous):

so sin theats = 5/8?

OpenStudy (accessdenied):

correct :)

OpenStudy (anonymous):

k then wha?

OpenStudy (accessdenied):

so from there, we just need to use inverse sine on both sides. calculating sin^(-1) (5/8) should give us the value of theta for -180 < theta < 180

OpenStudy (anonymous):

38.7?

OpenStudy (anonymous):

y u subtracting 180 tho?

OpenStudy (accessdenied):

woops, the range is -90 < theta < 90 for arcsin >.< keep mixing it up thinking about radians

OpenStudy (anonymous):

lol

OpenStudy (accessdenied):

but then recall this part: "obtuse angle" obtuse angles... how big is the obtuse angle from geometry?

OpenStudy (anonymous):

180?

OpenStudy (anonymous):

got it thnks!

OpenStudy (accessdenied):

sooo we're looking for an angle for \(\theta \) greater than 180 and less than 360. (definition of obtuse)

OpenStudy (anonymous):

thnks a lot! gn :D

OpenStudy (accessdenied):

basically, we found the angle 38.7. but thats much lower than 180. We need to use the angle 180+38.7 this time, because that angle lies in the correctinterval

OpenStudy (accessdenied):

gnight!